One number is three times another number. the sum of the number is 196. what are the numbers?
Answer by mananth(16949)

let one number be x
other number = 3x
..
3x+x=196
4x=196
/4
x= 196/4
x=49
...
the numbers are 49
& 3*49 = 147
m.ananth@hotmail.ca